OpenStudy (anonymous):

how would i graph the equation 2x+5y=10?

Nnesha (nnesha):

first you have to solve for y because in order to graph equation you need in slope intercept form which is y = mx+b

OpenStudy (anonymous):

okay what would be the next step?

Nnesha (nnesha):

next step to graph

Nnesha (nnesha):

y-intercept where line cross y -axis so first step to graph it ----> draw point (y-intercept) at y-axis

OpenStudy (anonymous):

2x -2x+5y=-2x+10 5/5y=5/-2x+10/5 y=-5/2x+2

OpenStudy (anonymous):

is that right? then I would use that answer to graph?

Nnesha (nnesha):

nope :(

Nnesha (nnesha):

 that what you got after subtracting by 2x now divide both side by 5